40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 15 12:28:03 2016, gautam, Update, General, Something has gone wrong - was there a power outage? Vac_15Feb2016.png
    Reply  Mon Feb 15 13:09:33 2016, Koji, Update, General, Something has gone wrong - was there a power outage? 
       Reply  Tue Feb 16 09:13:39 2016, Steve, Update, General, there was a power outage & EQ 20160215_power_outage.png
          Reply  Tue Feb 16 15:02:19 2016, ericq, Update, General, Power Glitch recovery 
             Reply  Tue Feb 16 23:42:22 2016, gautam, Update, General, Power Glitch recovery - arms recovered 
                Reply  Wed Feb 17 13:45:15 2016, ericq, Update, General, No monit on frontends 
             Reply  Thu Feb 25 11:02:36 2016, Steve, Update, General, Power Glitch again powerGlitch.png
                Reply  Thu Feb 25 11:32:04 2016, gautam, Update, General, Power Glitch again 
          Reply  Mon Feb 22 08:45:46 2016, Steve, Update, SUS, EQ 4.3 grabs ITMX-UL magnet EQ4.3LucerneV.pngITMX-UL.png
             Reply  Mon Feb 22 23:00:21 2016, ericq, Update, SUS, ITMX free, green is flashing 
                Reply  Tue Feb 23 09:51:34 2016, Steve, Update, SUS, ITMX free ITMX_FREED.pngITMX_FREE.png
                   Reply  Tue Feb 23 17:46:09 2016, gautam, Update, SUS, EQ  SUS_Summary_23Feb.png
             Reply  Tue Apr 5 08:55:51 2016, Steve, Update, SUS, local EQ 3.1m eq3.1Valencia.png
                Reply  Wed May 25 08:32:55 2016, Steve, Update, SUS, local EQ 3.5m 3.5Cucam.pnglocal3.5cucam.png
                   Reply  Fri Jun 3 14:24:13 2016, Steve, Update, SUS, local EQ 3.1 m eq3.1mMuscoyCa.png
                      Reply  Fri Jun 10 09:52:51 2016, Steve, Update, SUS, EQ 5.2 mag Borego Spring 5.2mBoregoSpring.pngeqBSs.png
                         Reply  Mon Jun 20 09:36:10 2016, Steve, Update, SUS, EQ 3.1 m Big Bear Lake, CA eq3.1BigBearLakeCA.png
Message ID: 11995     Entry time: Tue Feb 16 23:42:22 2016     In reply to: 11993     Reply to this: 11997
Author: gautam 
Type: Update 
Category: General 
Subject: Power Glitch recovery - arms recovered 

 I was able to realign the arms, lock them, and have run the dither align to maximize IR transmission - looks like things are back to normal now. For the Y-end, I used the green beam initially to do some coarse alignment of the ITM and ETM, till I was able to see IR flashes in the control room monitors. I then tweaked the alignment of the tip-tilts till I saw TEM00 flashes, and then enabled LSC. Once the arm was locked, I ran the dither align. I then tweaked ITMX alignment till I saw IR flashes in the X arm as well, and was able to lock it with minimal tweaking of ETMX. The LSC actuation was set to ETMX when the models were restarted - I changed this to ITMX actuation, and now both arms are locked with nominal IR transmissions. I will center all the Oplev spots tomorrow before I start work on getting the X green back - I've left the ETM Oplev servos on for now.

While I was working, I noticed that frame builder was periodically crashing. I had to run mxstream restart a few times in order to get CDS back to the nominal state. I wonder if this is a persistent effect of the date/time issues we were seeing earlier today?

ELOG V3.1.3-